Gage Corporation has two operating divisions in a semiautonomous organizational structure. Adams Division, located in the United States, produces a specialized electrical component that is an input to Bute Division, located in the south of England. Adams uses idle capacity to produce the component, which has a domestic market price of $13. Its variable costs are $4 per unit. Gage’s U.S. tax rate is 40 percent of income. In addition to the transfer price for each component received from Adams, Bute pays a $2 per unit shipping fee. The component becomes a part of its assembled product, which costs an additional $2 to produce and sells for an equivalent of $21. Bute could purchase the component from a Manchester (England) supplier for $11 per unit. Gage’s English tax rate is 70 percent of income. Assume that English tax laws permit transferring at either variable cost or market price. Required: a-1. What are the respective profits after tax for both the Adams Division and Bute Division of Gage if the transfer price is $4. a-2. What are the respective profits after tax for both the Adams Division and Bute Division of Gage if the transfer price is $13. a-3. What transfer price is economically optimal for Gage Corporation?
Answer #1

Answer a-1:

Adams Division:

Variable cost per unit = $4

Bute Division:

Input cost based on transfer price

Shipping cost = $2

Additional production and selling cost = $2 per unit

As such:

Adams Division:

Profit after tax per unit = (Transfer price - Variable cost) * (1 - Tax rate) = ($4 - $4) * (1 - 40%) = $0

Bute Division:

Profit after tax per unit = (Sale price - transfer price - Shipping cost - Additional production and selling cost ) * (1 - Tax rate)

= ($21 - $4 - $2 - $2) * (1 - 70%) = $3.90

Adams Division Bute Division $3.90 Profit after tax Per unit

For Gage corporation profit after tax per unit will be = $0 + $3.90 = $3.90

Answer a-2:

If the transfer price is $13:

Adams Division:

Profit after tax per unit = (Transfer price - Variable cost) * (1 - Tax rate) = ($13 - $4) * (1 - 40%) = $5.40

Bute Division:

Profit after tax per unit = (Sale price - transfer price - Shipping cost - Additional production and selling cost ) * (1 - Tax rate)

= ($21 - $13 - $2 - $2) * (1 - 70%) = $1.20

Adams Division Bute Division $5.40 $1.20 Profit after tax Per unit

For Gage corporation profit after tax per unit will be = $5.40 + $1.20 = $6.60

Answer a-3:

Since income tax is higher at 70% in UK (for Bute Division) and 40% in US (for Adam division), optimal transfer price from Adam to Bute will be the one, which is highest possible.

Given that English tax laws permit transferring at either variable cost or market price.

Hence:

The optimal transfer price is $13.

As we observe, for Gage corporation, when transfer price is $4, the profit after tax per unit = $3.90 but when transfer price is $13, for Gage Corporation profit after tax per unit is higher at $6.60

The optimal transfer price is $13.